# 3D-Aware Instance Segmentation and Tracking in Egocentric Videos

### v1

@@ 1 Introduction
Egocentric videos, which capture the world from a first-person perspective, are a focus of increasing attention in computer vision due to their importance in applications such as augmented reality and robotics. Among various tools for video analysis, object tracking is of particular importance, but also faces significant challenges, in the egocentric case. Most video object segmentation (VOS) methods [6, 20, 24, 37], in fact, assume that the videos contains slow, steady camera motions that keep the view centered on the object of interest [4, 27, 1]. In comparison, egocentric videos are taken from a first-person perspective, where the camera wearer’s movements introduce rapid and unpredictable changes in viewpoint. Additionally, objects frequently move in and out of the field of view, and thus are often partially or wholly occluded and/or truncated.
For example, in the EPIC KITCHENS dataset [10], the person taking the video might move a pan on top of a hob and leave it there for several minutes while moving around in the kitchen. During that time, they might observe more objects that look similar to the pan, which may cause an algorithm to incorrectly associate them to the pan itself. In general, video segmenters tend to lose track of the object partially or entirely due to occlusion or truncation. These issues are exacerbated when tracking multiple objects simultaneously.
@@
In order to address these shortcomings, we can look at how humans locate objects. An important cue that helps correct reassociation is object permanence, a concept that human infants develop very early [31]. Permanence captures the idea that objects do not cease to exist when they are not visible. Combined with spatial awareness, this means that the 3D location of objects at rest should not change when they are out of view or occluded.
@@ 2 Related Work
Point tracking-based methods have been pivotal in advancing VOS by providing a means to establish correspondences across frames. Many powerful point trackers have been recently proposed such as TAP-Vid [11] benchmark that focused on tracking physical points in a video and works such as CoTracker [16] and PIP [14]. CenterTrack [45] combined object detection with point tracking, leveraging the strengths of both approaches. TAPIR [12] trains an initial matching network (analogous to Seqformer) and an iterative refinement network (which focuses on continuous adjustments to predicted points’ positions), using synthetic data, to predict accurate point tracks. SAM-PT [30] is a point-centric interactive video segmentation, which propagates a sparse set of points, chosen by a user, to other frames.
@@
Exploiting 3D information in egocentric videos has been less explored due to the challenges of reconstructing dynamic objects. Following [2], EgoLifter [13] uses contrastive learning to lift 2D segmentations to 3D, while also using a transient prediction network to handle dynamic objects. Plizzari et al. [28] focus specifically on 3D tracking of dynamic objects, rather than segmenting or reconstructing them. They form 3D centroid tracks by lifting 2D centroids to 3D and matching observations based on 3D distance and visual similarity.
We build on this by integrating scene geometry and per-frame depth estimates to refine the associations of instance segmentations. We incorporate instance and category information from a base VOS model into our cost formulation, creating a more robust 3D-aware object tracking system that achieves long-term consistent object segmentations in egocentric videos.
@@ 3.1 Problem statement
We begin with an egocentric video sequence consisting of $N$ frames $\mathbf{I}^{t}$, $t\in\{1,...,N\}$, along with the output of an off-the-shelf 2D VOS model. This initial output contains a set of object tracks that, while partially correct, often contain errors – particularly when objects temporarily leave the field of view or are occluded. Our goal is to refine and reassemble these tracks, leveraging 3D information to correct errors and achieve more consistent long-term tracking. Crucially, we don’t discard the initial track IDs obtained from the 2D-only VOS model. Instead, we incorporate this information into our refinement process, using it as a valuable prior for maintaining object identities. The objective of the method is to compute a set of tracks for the entire video $\{\mathbf{T}^{N}_{i}\}$ with associated segment IDs $\{\tilde{s}_{i}^{N}\}$ that have the desired temporal consistency.
@@
Here, $\mathbf{1}(s_{i}^{t}=\tilde{s}_{j}^{t-1})$ is an indicator function. $\delta_{p}(b_{i,p}^{t},\tilde{b}_{j,p}^{t+1})$ is the consistency cost for the $p$-th attribute between segment $m_{i}^{t}$ in frame $t$ and track $T_{j}^{t-1}$. Importantly, one of these $\delta_{p}$ functions specifically accounts for the initial track IDs (c.f. Eq. 8), encouraging our optimization to maintain these associations when appropriate.
We use the Hungarian algorithm to solve for the new segment IDs and update the initial segment IDs only if optimisation cost from Eq. 2 is below a cost threshold $\gamma$. This ensures that our algorithm does not change associations when cost is too high. Notably, for new observations that don’t match any existing track (i.e., their matching cost exceeds $\gamma$), we initialize new tracks. Importantly, we don’t terminate tracks that fail to match with a new observation in the current frame. Instead, we maintain these tracks in our database, propagating their attributes from time $t-1$ to time $t$. This approach allows our method to handle temporary occlusions or brief disappearances of objects, maintaining object identity over longer periods.
@@ 3.4 Cost functions
The attributes used for refining the tracks are thus $\mathbf{b}_{i}^{t}=(l_{i}^{t},v_{i}^{t},\bar{c}_{i}^{t},\bar{s}_{i}^{t})$, consisting of the 3D location, the visual features, initial category label and initial instance label for the segment $m_{i}^{t}$ of frame $t$. Now, we define the cost functions $\delta_{p}$ used in Eq. 2 for these individual attributes.
@@
    Similar to [28], we model the location cost $\delta_{l}$ with the exponential distribution as follows:
@@
We refer the reader to Appendix 0.A for the implementation details and to Appendix 0.C for hyperparameter settings.
@@ 4.1 Benchmark and baselines
We evaluate our proposed method on 20 challenging scenes from the EPIC Fields [34] dataset. EPIC Fields comprises of complex real-world videos with a high diversity of activities and object interactions, making it an ideal testbed for our evaluation. The scenes we selected include varied lighting conditions, occlusions as well as objects intermittently being out of sight. Each video containing several thousand frames, averaging around 10 minutes in length.
@@ 4.2 Metrics
$\text{HOTA}=\frac{1}{|S|}\sum_{\alpha\in S}\text{HOTA}(\alpha)=\frac{1}{|S|}\sum_{\alpha\in S}\sqrt{\text{DetA}(\alpha)\times\text{AssA}(\alpha)}$
@@
$\text{DetA}(\alpha)=\frac{|\text{TP}_{\alpha}|}{|\text{TP}_{\alpha}|+|\text{FP}_{\alpha}|+|\text{FN}_{\alpha}|}$
@@
Whereas, AssA measures the tracker’s ability to maintain consistent object identities over time:
$\text{AssA}(\alpha)=\frac{1}{|\text{TP}_{\alpha}|}\sum_{c\in\text{TP}_{\alpha}}\frac{|\text{TPA}(c)|}{|\text{TPA}(c)|+|\text{FPA}(c)|+|\text{FNA}(c)|}$
@@ 4.4 Ablations
Sensitivity to hyperparameters.
We evaluate the sensitivity of our method by varying the values of the 4 hyperparameters: $\alpha_{s}$, $\alpha_{c}$, $\alpha_{l}$, $\alpha_{v}$ in the cost-matching formulation Eq. 2 from Eqs. 7, 5, 6 and 8. We perform this analysis on a subset of 5 videos, using 3 representative values for each hyperparameter, resulting in $813^{4}=$ 81 configurations. Figure 3 shows that $57$ out of these $81$ hyperparameter configurations lead to a HOTA score in the range $27.2\pm 0.2$, which shows the robustness of our method to these parameters. There are some configurations, e.g. when $\alpha_{s}=100$ or $\alpha_{c}=100$, that lead to a degradation in performance.
@@
4.5 Downstream applications
Reconstruction of objects.
Amodal segmentation.
@@ 5 Conclusion
In this paper, we presented a novel 3D-aware approach to instance segmentation and tracking in egocentric videos, addressing the unique challenges of first-person perspectives. By integrating 3D information, our method significantly improves tracking accuracy and segmentation consistency compared to state-of-the-art 2D approaches, especially over long periods. Our ablation studies highlight the importance of 3D information and the category as well as instance cost terms in matching, while also showing robustness to hyperparameter changes. Beyond improved tracking, our approach enables valuable downstream applications such as high-quality 3D object reconstructions and amodal segmentation. This work demonstrates the power of incorporating 3D awareness into egocentric video analysis, opening up new possibilities for robust object tracking in challenging first-person scenarios.
Acknowledgements
We are grateful for funding from EPSRC AIMS CDT EP/S024050/1 and AWS (Y. Bhalgat), NAVER LABS Europe (V. Tschernezki), ERC-CoG UNION 101001212 (A. Vedaldi and I. Laina), EPSRC VisualAI EP/T028572/1 (I. Laina, A. Vedaldi and A. Zisserman), and Royal Academy of Engineering RF\201819\18\163 (J. Henriques). We thank Chiara Plizzari for sharing implementation details of the OSNOM baseline, and Ahmad Dar Khalil for helping with the annotations for our evaluation.
@@ Appendix 0.A Implementation details
In the EPIC Fields [34] dataset, the per-frame camera pose and intrinsics are obtained using COLMAP [32], which also provides a sparse point cloud representing the static parts of the scene. We follow [28], and obtain the depth maps, $\mathbf{D}_{t}$ in Eq. 3, by first calculating a mesh from the sparse point clouds, and then aligning the predictions of a pre-trained depth model to the mesh through shift and scale transformations. We use Depth Anything [38], a state-of-the-art monocular depth estimation model. We obtain the scale-shift parameters for each depth map by optimizing the L1 distance between the transformed depth map and the rasterized mesh depth. We use the DINOv2 [26] encoder to compute the visual features for segments as described in Equation 4. When optimizing the tracking cost, the visual feature for a “track” is computed as the average of the visual features of the most recent $100$ observations assigned to the track.
Appendix 0.B Details on frame-by-frame Tracking Cost optimization
@@ Appendix 0.D Evaluation Data
Appendix 0.E Comparison with OSNOM [28] for Video Segmentation
Our approach is inspired by OSNOM [28]. But while OSNOM focuses only on coarse and class-agnostic 3D centroid tracking, ours is tailored to be more suitable for long-term video instance segmentation by taking in account the initial categorical information provided by the base 2D model. To facilitate a direct comparison, we simulate OSNOM’s functionality within our framework by disabling the instance and category terms in our cost formulation, relying solely on 3D location and visual feature costs.
Table 4: Comparison of our method with OSNOM [28] for egocentric video instance segmentation. *OSNOM is simulated within our framework by disabling the instance and category terms in our cost formulation.
@@ Appendix 0.E Comparison with OSNOM [28] for Video Segmentation
OSNOM (w/ DEVA)* 11.31 6.49 24.16 12.98
Ours (w/ DEVA) 27.72 18.38 43.90 26.63
As shown in Table 4, this OSNOM-like configuration results in a significant performance drop, with metrics falling below even the initial performance obtained using the base 2D model, DEVA [6].
This is due to two reasons. First, without the “instance” cost, the model completely ignores the initial tracks provided by DEVA. Second, without the “category” cost, the model often confuses objects across categories (e.g. pot vs sink, knife vs spoon). Since the Detection Accuracy (DetA) for a video/scene is computed on predicted instances per class, this leads to a severely low DetA for OSNOM on account of various misclassified instances. This comparison underscores the importance of our additional cost terms in maintaining robust long-term tracking performance in egocentric settings.
Appendix 0.F Details on Obtaining Amodal Segmentations
As briefly described in Section 4.6 of the main paper, our method for obtaining amodal segmentations builds upon the long-term instance segmentations and coarsely tracked 3D centroids generated by our primary algorithm. This section elaborates on the process, which involves three main steps: identifying static object frames, 3D object reconstruction, and amodal segmentation projection.

### v2

@@ 1 Introduction
Egocentric videos, which capture the world from a first-person perspective, are a focus of increasing attention in computer vision due to their importance in applications such as augmented reality and robotics. Among various tools for video analysis, object tracking is of particular importance, but also faces significant challenges, in the egocentric case. Most video object segmentation (VOS) methods [8, 23, 28, 43], in fact, assume that the videos contain slow, steady camera motions that keep the view centered on the object of interest [5, 31, 2]. In comparison, egocentric videos are taken from a first-person perspective, where the camera wearer’s movements introduce rapid and unpredictable changes in viewpoint. Additionally, objects frequently move in and out of the field of view, and thus are often partially or wholly occluded and/or truncated.
For example, in the EPIC KITCHENS dataset [12], the person recording the video might move a pan on top of a hob and leave it there for several minutes while moving around in the kitchen. During that time, they might observe more objects that look similar to the pan, which may cause an algorithm to incorrectly associate them to the pan itself. In general, video segmenters tend to lose track of the object partially or entirely due to occlusion or truncation. These issues are exacerbated when tracking multiple objects simultaneously.
@@
In order to address these shortcomings, we can look at how humans locate objects. An important cue that helps correct reassociation is object permanence, a concept that human infants develop very early [36]. Permanence captures the idea that objects do not cease to exist when they are not visible. Combined with spatial awareness, this means that the 3D location of objects at rest should not change when they are out of view or occluded. It has previously been explored for egocentric videos in ‘Out of Sight, Not Out of Mind’ (OSNOM) [32].
@@ 2 Related Work
Point tracking-based methods have been pivotal in advancing VOS by providing a means to establish correspondences across frames. Many powerful point trackers have been recently proposed such as TAP-Vid [13] benchmark that focused on tracking physical points in a video and works such as CoTracker [19] and PIP [17]. CenterTrack [52] combined object detection with point tracking, leveraging the strengths of both approaches. TAPIR [14] trains an initial matching network (analogous to SeqFormer) and an iterative refinement network (which focuses on continuous adjustments to predicted points’ positions), using synthetic data, to predict accurate point tracks. SAM-PT [35] is a point-centric interactive video segmentation method, which propagates a sparse set of points, chosen by a user, to other frames.
@@
Exploiting 3D information in egocentric videos has been less explored due to the challenges of reconstructing dynamic objects. Following [3], EgoLifter [16] uses contrastive learning to lift 2D segmentations to 3D, while also using a transient prediction network to handle dynamic objects. Plizzari et al. [32] focus specifically on 3D tracking of dynamic objects, rather than segmenting or reconstructing them. They form 3D centroid tracks by lifting 2D centroids to 3D and matching observations based on 3D distance and visual similarity. We follow [32], in that we lift objects to 3D using estimated depth, and initialise, match and update tracks based on 3D location and DINOv2 [30] feature similarity. However, we also incorporate instance and category information from a base VOS model into our cost formulation, creating a more robust 3D-aware object tracking system that excels in refining imperfect or noisy input 2D object tracks, achieving superior long-term object consistency as compared to existing 2D tracking methods.
@@ 3.1 Problem statement
We begin with an egocentric video sequence consisting of $N$ frames $\mathbf{I}^{t}$, $t\in\{1,\dots,N\}$, along with the output of an off-the-shelf 2D VOS model. The objective of the method is to compute a set of tracks for the entire video $\{\mathbf{T}^{N}_{i}\}$ with associated segment IDs $\{\tilde{s}_{i}^{N}\}$ that have the desired temporal consistency. The initial output contains a set of object tracks that, while partially correct, often contain errors — particularly when objects temporarily leave the field of view or are occluded. Our goal is to refine and reassemble these tracks, leveraging 3D information to correct errors and achieve more consistent long-term tracking. Crucially, we don’t discard the initial track IDs obtained from the 2D-only VOS model. Instead, we incorporate this information into our refinement process, using it as a valuable prior for maintaining object identities. In this manner, we go beyond the previous 3D aware matching, initialisation and matching method [32] that we build upon.
@@
Here, $\mathbf{1}(s_{i}^{t}=\tilde{s}_{j}^{t-1})$ is an indicator function. $\delta_{p}(b_{i,p}^{t},\tilde{b}_{j,p}^{t-1})$ is the consistency cost for the $p$-th attribute between segment $m_{i}^{t}$ in frame $t$ and track $T_{j}^{t-1}$. Importantly, one of these $\delta_{p}$ functions specifically accounts for the initial track IDs (c.f. Eq. 8), encouraging our optimization to maintain these associations when appropriate.
We use the Hungarian algorithm to solve for the new segment IDs and update the initial segment IDs only if the optimisation cost from Eq. 2 is below a cost threshold $\gamma$. This ensures that our algorithm does not change associations when the cost is too high. Notably, for new observations that don’t match any existing track (i.e., their matching cost exceeds $\gamma$), we initialize new tracks. Importantly, we do not terminate tracks that fail to match with a new observation in the current frame. Instead, we maintain these tracks in our database, propagating their attributes from time $t-1$ to time $t$. This approach allows our method to handle temporary occlusions or brief disappearances of objects, maintaining object identity over longer periods.
@@ 3.4 Cost functions
The attributes used for refining the tracks are thus $\mathbf{b}_{i}^{t}=(l_{i}^{t},v_{i}^{t},\bar{c}_{i}^{t},\bar{s}_{i}^{t})$, consisting of the 3D location, the visual features, initial category label and initial instance label for the segment $m_{i}^{t}$ of frame $t$. Now, we define the cost functions $\delta_{p}$ used in Eq. 2 for these individual attributes. We follow [34, 32] for the first two:
@@
    We model the 3D location cost $\delta_{l}$ with the exponential distribution as follows:
@@
We refer the reader to Appendix 0.A for the implementation details and to Section 0.A.2 for hyperparameter settings.
@@ 4.1 Benchmark and baselines
We evaluate our proposed method on 20 challenging scenes from the EPIC Fields [40] dataset. EPIC Fields comprises of complex real-world videos with a high diversity of activities and object interactions, making it an ideal testbed for our evaluation. The selected videos include varied lighting conditions, occlusions, objects that disappear from sight, and have an average length of 10 minutes. To further demonstrate our method’s capability, we also evaluate it on the Ego4D [15] dataset and report the results in Table 5.
@@
In Appendix 0.E, we carry out an ablation disabling the category and instance terms in our cost function. This brings our approach somewhat closer to the method of OSNOM [32].
@@ 4.2 Metrics
$\text{HOTA}=\frac{1}{|S|}\sum_{\alpha\in S}\text{HOTA}(\alpha)=\frac{1}{|S|}\sum_{\alpha\in S}\sqrt{\text{DetA}(\alpha)\times\operatorname{AssA}(\alpha)}$
@@
$\text{DetA}(\alpha)=\frac{|\mathrm{TP}_{\alpha}|}{|\mathrm{TP}_{\alpha}|+|\text{FP}_{\alpha}|+|\text{FN}_{\alpha}|}$
@@
AssA measures the tracker’s ability to maintain consistent object identities over time:
$\operatorname{AssA}(\alpha)=\frac{1}{|\mathrm{TP}_{\alpha}|}\sum_{c\in\mathrm{TP}_{\alpha}}\frac{\operatorname{TPA}(c)|}{|\operatorname{TPA}(c)|+|\operatorname{FPA}(c)|+|\operatorname{FNA}(c)|}$
@@ 4.4 Ablations
Comparison with other plug-and-play tracking methods
The above results demonstrated our method’s generalization capability by combining with with two state-of-the-art methods, DEVA [8] and MASA [23]. To further highlight our method’s versatility, we compare it with other existing plug-and-play tracking algorithms, namely BoTSORT [1], ByteTrack [49], OCSORT [6] and DeepOCSORT [26]. We use the same ReID model with all 4 tracking methods and use OWLv2 [27] as the 2D segmentation model for fair comparisons. Table 3 shows that all four methods perform less favourably than DEVA [8] even while using the same base 2D model, and thus are outperformed by our method which further refines the tracks from DEVA.
Table 3: Comparison with other plug-and-play tracking methods.
Method HOTA DetA AssA IDF1
BoTSORT [1] 12.83 7.81 24.23 10.30
ByteTrack [49] 20.08 16.31 27.56 16.94
OCSORT [6] 21.90 17.94 29.28 18.95
DeepOCSORT [26] 22.63 17.98 31.31 19.88
DEVA [8] 25.14 18.40 36.72 22.17
Ours (w/ DEVA) 27.72 18.38 43.90 26.63
@@ 5 Conclusion
In this paper, we presented a novel 3D-aware approach to instance segmentation and tracking in egocentric videos, addressing the unique challenges of first-person perspectives. By integrating 3D information, our method significantly improves tracking accuracy and segmentation consistency compared to state-of-the-art 2D approaches, especially over long periods. Our ablation studies highlight the importance of 3D information and the category as well as instance cost terms in matching, while also showing robustness to hyperparameter changes. Beyond improved tracking, our approach enables valuable downstream applications such as high-quality 3D object reconstructions and amodal segmentation. This work demonstrates the power of incorporating 3D awareness into egocentric video analysis, opening up new possibilities for robust object tracking in challenging first-person scenarios. {credits}
5.0.1 Acknowledgements
We’re funded by EPSRC AIMS CDT EP/S024050/1 and AWS (Y. Bhalgat), NAVER LABS Europe (V. Tschernezki), ERC-CoG UNION 101001212 (A. Vedaldi and I. Laina), EPSRC VisualAI EP/T028572/1 (I. Laina, A. Vedaldi and A. Zisserman), and Royal Academy of Engineering RF\201819\18\163 (J. Henriques). We thank Chiara Plizzari for sharing evaluation details of the OSNOM baseline, and Ahmad Dar Khalil for helping with the annotations for our evaluation.
@@ Appendix 0.A Implementation details
In the EPIC Fields [40] dataset, the per-frame camera pose and intrinsics are obtained using COLMAP [37], which also provides a sparse point cloud representing the static parts of the scene. We follow [32], and obtain the depth maps, $\mathbf{D}_{t}$, by first calculating a mesh from the sparse point clouds, and then aligning the predictions of a pre-trained depth model to the mesh through shift and scale transformations. We use Depth Anything [44], a state-of-the-art monocular depth estimation model. We obtain the scale-shift parameters for each depth map by optimizing the L1 distance between the transformed depth map and the rasterized mesh depth. We use the DINOv2 [30] encoder to compute the visual features for segments. When optimizing the tracking cost, the visual feature for a “track” is computed as the average of the visual features of the most recent $100$ observations assigned to the track.
0.A.1 Details on frame-by-frame Tracking Cost optimization
@@ 0.A.3 Evaluation Data
Appendix 0.B Additional results on the Ego4D dataset
To further demonstrate our method’s applicability, we also include results on a few select scenes from the Ego4D [15] dataset. We follow the same evaluation protocol used with the EPIC Fields dataset and utilize Egotracks [39] for the ground-truth track annotations.
Table 5 shows that our method can consistently refine the tracks obtained by DEVA [8] on 3 videos. We use the same hyperparameters for these experiments as the ones used with the EPIC Fields datasets described in Sec. 0.A.2. The video lengths are 2-3$\times$ shorter compared to EPIC Fields, which we believe results in smaller margins of improvement as compared to the ones showed in the main paper.
Table 5: Results on the Ego4D [15] dataset.
Video ID DEVA [8] Ours (w/ DEVA)
HOTA DetA AssA IDF1 HOTA DetA AssA IDF1
8b47ac19-7c4f-47d2- b5d0-755b524b66b2 15.29 12.26 22.74 13.56 17.40 12.22 28.15 14.98
9f5253af-acc3-40ca- b8bf-7b931f875bd7 12.37 9.25 18.61 10.43 14.38 9.25 23.04 12.33
bff3d583-ca3b-44b8- 9740-3b34c5a8d7a9 21.58 13.60 36.21 18.27 23.73 13.58 43.96 20.16
Appendix 0.C On-device Inference Runtime Analysis
Our method processes egocentric videos in an online manner. While we compute meshes in advance, we track the objects online with the entire pipeline running at 20FPS on a A6000 GPU. Each of our method’s components runs as follows: DINOv2 at 43FPS, the lifting to 3D with DepthAnythingV2 at 23FPS, the prediction of segmentation masks with OWLv2 (run every 5 frames) at 31FPS, the temporal propagation with DEVA [8] at 25 FPS. We use torch.cuda.stream for asynchronous execution of all models on the same GPU.
Appendix 0.D Sensitivity to Hyperparameters
We evaluate the sensitivity of our method by varying the values of the 4 hyperparameters: $\alpha_{s}$, $\alpha_{c}$, $\alpha_{l}$, $\alpha_{v}$ in the cost-matching formulation. We perform this analysis on a subset of 5 videos, using 3 representative values for each hyperparameter, resulting in $813^{4}=$ 81 configurations. Figure 4 shows that $57$ out of these $81$ hyperparameter configurations lead to a HOTA score in the range $27.2\pm 0.2$, which shows the robustness of our method to these parameters. There are some configurations, e.g. when $\alpha_{s}=100$ or $\alpha_{c}=100$, that lead to a degradation in performance.
Appendix 0.E Ablation without category and instance terms
Here, we ablate our method by disabling the instance and category terms in the cost formulation, relying solely on 3D location and visual feature costs. Note that this ablation is compared with the “full” version of our method on the object tracking task.
Table 6: Ablation of our method without category and instance terms in the cost formulation, evaluated on the egocentric object tracking task.
@@ Appendix 0.E Ablation without category and instance terms
Ours (without cat and ins terms) 11.31 6.49 24.16 12.98
Ours (full version) 27.72 18.38 43.90 26.63
As shown in Table 6, this ablation results in a significant performance drop, with metrics falling below even the initial performance obtained using the base 2D model, DEVA [8]. This is due to two reasons. First, without the “instance” cost, the model completely ignores the initial tracks provided by DEVA. Second, without the “category” cost, the model often confuses objects across categories (e.g. pot vs sink, knife vs spoon). Since the Detection Accuracy (DetA) for a video/scene is computed on predicted instances per class, this leads to a severely low DetA on account of various misclassified instances. This comparison underscores the importance of our additional cost terms in maintaining robust long-term tracking performance in egocentric settings.
Note, the ablation brings the method somewhat closer to the the OSNOM-like configuration of [32]. Although our method was inspired by OSNOM, we can’t carry out a direct comparison as the authors of [32] have not yet released their code, or details of the videos used in their evaluation.
Appendix 0.F Downstream applications
Reconstruction of objects.
Amodal segmentation.
Appendix 0.G Details on Obtaining Amodal Segmentations
This section elaborates on the process of obtaining amodal segmentations, which involves three main steps: identifying static object frames, 3D object reconstruction, and amodal segmentation projection.